OpenStudy (australopithecus):

can someone show me how to use \[(1+x)^{k}=\sum_{n=0}^{\infty}\left(\begin{matrix}k \\ n\end{matrix}\right)x^n\] to convert the function: (2+x)^(-3) into a series?

OpenStudy (amistre64):

looks like a binomal thrm setup

OpenStudy (amistre64):

the longer method is just to carry out the division

OpenStudy (anonymous):

Let u=1+x, and then form a series in terms of u.

OpenStudy (amistre64):

\[\frac{1}{(2+x)^3}\] \[\frac{1}{8+12x+6x^2+x^3}\] 1/8 _____________________ 8+12x+6x^2+x^3 ) 1 -(1 + 3x/2 + 3x^2/4 +x^3/8) --------------------------- - 3x/2 - 3x^2/4 - x^3/8 etc .....

OpenStudy (amistre64):

1/8 - 3x/16 + 3x^2/32 - x^3/64 ____________________________ 8+12x+6x^2+x^3 ) 1 -(1 + 3x/2 + 3x^2/4 +x^3/8) --------------------------- - 3x/2 - 3x^2/4 - x^3/8 -(-3x/2 - 9x^2/4 - 9x^3/8 -3x^4/16) ----------------------------------- 3x^2/4 + 8x^3/8 +3x^4/16 -(3x^2/4 + 9x^3/8 + 9x^4/16+3x^5/32) ------------------------------------ -x^3/8 - 6x^4/16 - 3x^5/32 hmmm, i was hoping to see a pattern i could restructure :/

OpenStudy (australopithecus):

It states in my notes that I should use \[cn = \frac{f^{n}(0)}{n!}\]

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=1%2F%282%2Bx%29%5E2&lk=4 maybe, its something i havent played about with much tho so i cant verify. And it seems i can divide to well on a keyboard :)

OpenStudy (amistre64):

ugh, or type into the wolf that well either; i squared that other one .... hmm, looks like i was heading in the right direction for some of it tho :) http://www.wolframalpha.com/input/?i=1%2F%282%2Bx%29%5E3

OpenStudy (amistre64):

your notes indicate a taylor expansion

OpenStudy (amistre64):

the nth coeef is f^(n) (x=0)/n!

OpenStudy (australopithecus):

right but I still have no idea how to convert these using the standard identity, I may just use another method if it comes up on my exam